# Dynamics of Moral Behavior in Heterogeneous Populations of Learning Agents

### v4

@@ Introduction
In this Markov game (Littman 1994), over time both players learn to make selections and take actions by observing past interactions with various opponents. Each agent uses Q-Learning (Watkins and Dayan 1992) to update the state-action value function (i.e., an estimate of the cumulative expected return over time). Given a large number of possible selection states, we approximate using a neural network. Each agent maintains two internal models (one for partner selection, and one for playing the dilemma) - for consistency, we use function approximation for learning both models, and each is parametrized by a fully connected network with a single hidden layer of size .
@@ Appendix B Cooperation by Each Player Type across Populations
![[Uncaptioned image]](extracted/5757219/plots/appendix/manypopulaitons_actions_aggSelfish_avgacrossruns.png)
![[Uncaptioned image]](extracted/5757219/plots/appendix/manypopulaitons_actions_aggUtilitarian_avgacrossruns.png)
![[Uncaptioned image]](extracted/5757219/plots/appendix/manypopulaitons_actions_agganti-Utilitarian_avgacrossruns.png)
![[Uncaptioned image]](extracted/5757219/plots/appendix/manypopulaitons_actions_aggDeontological_avgacrossruns.png)
![[Uncaptioned image]](extracted/5757219/plots/appendix/manypopulaitons_actions_aggmalicious_Deontological_avgacrossruns.png)
@@
![[Uncaptioned image]](extracted/5757219/plots/appendix/manypopulaitons_actions_aggVirtueEthics_equality_avgacrossruns.png)
![[Uncaptioned image]](extracted/5757219/plots/appendix/manypopulaitons_actions_aggVirtueEthics_inequality_avgacrossruns.png)
![[Uncaptioned image]](extracted/5757219/plots/appendix/manypopulaitons_actions_aggVirtueEthics_kindness_avgacrossruns.png)
![[Uncaptioned image]](extracted/5757219/plots/appendix/manypopulaitons_actions_aggVirtueEthics_aggressiveness_avgacrossruns.png)
@@ Appendix C Selection Patterns between Player Types
In addition to the analysis of selections made over the entire simulation in the two populations presented in the main body of the paper (majority-S and majority-De, see Figure 4a & 4b), we show selections made by each individual player in every population in Figures 8a & 8b (as a heat map) and Figures 9a & 9b (as networks involving the top 10% of the pairs in terms of number of interactions). The latter show the most common patterns of interaction among players in these artificial societies - for example, in the majority-De population, the De players select other types of opponents more often than each other.
@@
Figure 9a: Selections made by individual players in populations majority-S, majority-Ut, majority-aUt, majority-De and majority-mDe. Edges are weighted by the number of selections made between the pair of players. We visualize the most common interactions, i.e., above the 90th percentile, over all 30000 episodes. More popular players are placed more centrally using a force-directed algorithm.
@@
Figure 9b: Selections made by individual players in populations majority-V-Eq, majority-V-In, majority-V-Ki and majority-V-Ag. Edges are weighted by the number of selections made between the pair of players. We visualize the most common interactions, i.e., above the 90th percentile, over all 30000 episodes. More popular players are placed more centrally using a force-directed algorithm.

### v5

@@ Introduction
In this Markov game (Littman 1994), over time both players learn to make selections and take actions by observing past interactions with various opponents. Each agent uses Q-Learning (Watkins and Dayan 1992) to update the state-action value function (i.e., an estimate of the cumulative expected return over time). Given a large number of possible selection states, we approximate using a neural network. Each agent maintains two internal models (one for partner selection, and one for playing the dilemma) - for consistency, we use function approximation for learning both models, and each is parameterized by a fully connected network with a single hidden layer of size .
@@ Appendix B Cooperation by Each Player Type across Populations
![[Uncaptioned image]](extracted/5778955/plots/appendix/manypopulaitons_actions_aggSelfish_avgacrossruns.png)
![[Uncaptioned image]](extracted/5778955/plots/appendix/manypopulaitons_actions_aggUtilitarian_avgacrossruns.png)
![[Uncaptioned image]](extracted/5778955/plots/appendix/manypopulaitons_actions_agganti-Utilitarian_avgacrossruns.png)
![[Uncaptioned image]](extracted/5778955/plots/appendix/manypopulaitons_actions_aggDeontological_avgacrossruns.png)
![[Uncaptioned image]](extracted/5778955/plots/appendix/manypopulaitons_actions_aggmalicious_Deontological_avgacrossruns.png)
@@
![[Uncaptioned image]](extracted/5778955/plots/appendix/manypopulaitons_actions_aggVirtueEthics_equality_avgacrossruns.png)
![[Uncaptioned image]](extracted/5778955/plots/appendix/manypopulaitons_actions_aggVirtueEthics_inequality_avgacrossruns.png)
![[Uncaptioned image]](extracted/5778955/plots/appendix/manypopulaitons_actions_aggVirtueEthics_kindness_avgacrossruns.png)
![[Uncaptioned image]](extracted/5778955/plots/appendix/manypopulaitons_actions_aggVirtueEthics_aggressiveness_avgacrossruns.png)
@@ Appendix C Selection Patterns between Player Types
In addition to the analysis of selections made over the entire simulation in the two populations presented in the main body of the paper (majority-S and majority-De, see Figure 4a & 4b), we show selections made by each individual player in every population in Figures 8a & 8b (as a heat map) and Figures 9a & 9b (as networks involving the top 15% of the pairs in terms of number of interactions observed). The latter show the most common patterns of interaction among players in these artificial societies - for example, in the majority-De population, the De players select other types of opponents more often than each other.
@@
Figure 9a: Selections made by individual players in populations majority-S, majority-Ut, majority-aUt, majority-De and majority-mDe. Edges are weighted by the number of selections made between the pair of players. We visualize the most common interactions, i.e., above the 85th percentile, over all 30000 episodes. More popular players are placed more centrally using a force-directed algorithm.
@@
Figure 9b: Selections made by individual players in populations majority-V-Eq, majority-V-In, majority-V-Ki and majority-V-Ag. Edges are weighted by the number of selections made between the pair of players. We visualize the most common interactions, i.e., above the 85th percentile, over all 30000 episodes. More popular players are placed more centrally using a force-directed algorithm.